OpenStudy (anonymous):

will fan and medal!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! 1. A model rocket is launched from the ground with an initial velocity of 160 ft/sec. a. How long will it take the rocket to reach its maximum height? Show all work in the space provided.

OpenStudy (anonymous):

you can use one of the equations of motion for constant acceleration. In this case the acceleration is due to gravity and = 32 ft /s^2 v = u + at at maximum height v = final velocity = 0 , u = 160 and a = -32 ( as the acceleration acts in opposite direction to motion) 0 = 160 - 32t t = 160/32 secnds

OpenStudy (anonymous):

b. Assume the model rocket’s parachute failed to deploy and the rocket fell back to the ground. How long would it take the rocket to return to Earth from the time it was launched? Show all work in the space provided.

OpenStudy (anonymous):

this is the whole thing

OpenStudy (anonymous):

what do u mean the question yea but I have other ones to answer to

OpenStudy (anonymous):

The rocket is on the ground at t+=+0 or t+=+10 seconds. The first time value is the starting point whereas the second time value is the ending point. So the rocket starts on the ground at t = 0 seconds, flies up, then comes back down and hits the ground at t = 10 seconds. The total flight time is 10 seconds.